How much energy would be required to heat 45 g of substance x from a liquid at 15 degrees C to a gas at 75 degrees C
1
Expert's answer
2020-04-15T01:51:33-0400

Solution.

"Q = c \times m \times \Delta t"

"Q = 4200 \times 45 \times (75-15) = 1134 \times 10^4 \ J"

Answer:

"Q = 1134 \times 10^4 \ J"